# IrrMap: A Large-Scale Comprehensive Dataset for Irrigation Method Mapping

### v1

@@ 8. Limitation of the dataset
The IrriMap dataset and trained model are currently limited to the western United States, which may restrict its applicability to regions with different climatic and agricultural conditions. All collected irrigation mapping data sources assume that irrigation practices and crop cultivation remain consistent throughout a season, which may not account for intra-seasonal variations due to environmental or management changes. Additionally, since the dataset is merged from multiple sources, it is subject to data inconsistencies, noise, and rasterization artifacts that may impact accuracy. Moreover, during the irrigation mapping, we drop the irrigated lands that have multiple or ambiguous irrigation labels. Furthermore, cropland data is generated through modeling, making it important not to over-rely on it for absolute ground truth. Groundwater data is not available in our IrrMap, limiting the model’s ability to assess subsurface water interactions.
@@ Acknowledgements.
This material is based upon work supported by the AI Research Institutes program supported by NSF and USDA-NIFA under the AI Institute: Agricultural AI for Transforming Workforce and Decision Support (AgAID) award No. 2021-67021-35344.
@@
Appendix A Properties of datasets
@@
Appendix B Data Analysis
(a) Arizona
(b) Colorado
(c) Utah
(d) Washington
Appendix C Evaluation Metrics
Appendix D Benchmarking
Additional Experiments. In Table 8, we present the overall performance on the training set. For training on Sentinel data, we use the same hyper-parameters and model architecture as Landsat. We randomly sample 100K training and test data from the IrrMap Sentinel dataset and train the model for five epochs.
The results indicate that the inclusion of additional bands, such as crop type and NDVI, has a notable impact on irrigation classification performance for both LandSat and Sentinel datasets. In LandSat, the RGB+CROP modality consistently outperforms other configurations, particularly in Flood and Sprinkler irrigation, suggesting that crop type information significantly enhances classification accuracy. Similarly, RGB+NDVI improves performance compared to standard RGB, but it does not surpass RGB+CROP, highlighting the greater influence of crop data over vegetation indices in LandSat imagery. For Sentinel, the impact of additional bands is even more pronounced, with RGB+CROP achieving the highest F1-scores and IoU across all irrigation methods, particularly for Sprinkler and Drip irrigation, where detailed crop information enhances model differentiation. RGB+LAND has also shown similar impact. The RGB+NDVI configuration also improves performance, especially in Sprinkler irrigation, reinforcing the value of vegetation indices at Sentinel’s higher resolution. Overall, the inclusion of auxiliary bands positively influences classification performance, with crop type data playing a critical role in both datasets.
Table 7. Hyperparameter details for supervised ESAWC experiments
Parameters
Task Type Irrigation Method Segmentation
Loss Function Cross Entropy
Input Dimensions $224\times 224$
Input Channels: RGB 3
Input Channels:RGB+CROP 24
Input Channels: RGB+LAND 4
Input Channels: RGB+NDVI 4
Output Dimensions $224\times 224$
Output Channels 4
Optimizer Adam
Learning Rate $1.00\times 10^{-2}$
No. Epochs (Landsat) 10
No. Epochs (Sentinel) 5
Model Selection Best (val)
Framework Pytorch Lightening
Table 8. Irrigation Classification Performance for Different Modalities, reporting Precision (P), Recall (R), F1-score (F), and Intersection over Union (IoU) for Flood, Sprinkler, and Drip irrigation methods across LandSat and Sentinel data.
Satellite Modality Flood Sprinkler Drip
P R F IoU P R F IoU P R F IoU

### v2

@@ 8. Limitation of the dataset
The IrriMap dataset and trained model are currently limited to the western United States, which may restrict its applicability to regions with different climatic and agricultural conditions. All collected irrigation mapping data sources assume that irrigation practices and crop cultivation remain consistent throughout a season, which may not account for intra-seasonal variations due to environmental or management changes. Since the dataset is merged from multiple sources, it is subject to data inconsistencies, noise, and rasterization artifacts that may impact accuracy. The crop mask relies on the cropland data, which is a model-generated mask. Any errors in model predictions can percolate into the classification tasks. Therefore, it is important to not over-rely on this dataset as well as substitute it with better equivalent datasets where available. Groundwater data is not available in our IrrMap, limiting the model’s ability to assess subsurface water interactions. While our mapping of original irrigation labels to three broad categories (drip, sprinkler, and flood) facilitates standardized analysis across diverse data sources, we acknowledge that this abstraction may obscure nuanced differences between specific technologies (e.g., center pivot vs. traveling gun). This trade-off was necessary to enable scalable, multi-state analysis, but future iterations of IrrMap will aim to preserve finer-grained irrigation method distinctions to support precision water management research.
@@ Acknowledgements.
This work was supported in part by the NSF and USDA-NIFA under the AI Institute: Agricultural AI for Transforming Workforce and Decision Support (AgAID) award No. 2021-67021-35344, and by University of Virginia Strategic Investment Fund award number SIF160.
@@
More details are available in our online Appendix (Mandal et al., 2025).
Appendix A Evaluation Metrics
Appendix B Benchmarking
Table 6. Hyperparameter Details
Data Setup
Task Type Input Dim RGB RGB+CROP RGB+LAND RGB+NDVI Output Ch.
Irrigation Method Segmentation $224\times 224$ 3 24 4 4 4
Training Setup
Loss Function Optimizer Learning Rate Epochs (L) Epochs (S) Model Sel. Framework
Cross Entropy Adam $1\times 10^{-2}$ 10 5 Best (val) PyTorch Lightning
Appendix C Data Analysis
(a) Arizona
(b) Colorado
(c) Utah
(d) Washington
Appendix D Properties of datasets